OpenStudy (anonymous):

find the slope and y intercept of the line 9x +2y=31

OpenStudy (anonymous):

slope is -9/2

OpenStudy (anonymous):

Change into Slope-Intercept form. y = mx + b m is slope b is y-intercept 9x + 2y = 31 Subtract 9x. 2y = -9x + 31 Divide both sides by 2. y = -9/2x + 31/2 or as decimal: y = -4.5x + 15.5 What's the slope? What's the y-intercept?

OpenStudy (anonymous):

thank you both

OpenStudy (anonymous):

welcome. (:

OpenStudy (anonymous):

ok